# e-Profits: A Business-Aligned Evaluation Metric for Profit-Sensitive Customer Churn Prediction

### v1

@@ 3.4 Retention Rate Estimation via Survival Analysis
  • Average Retention Rate (ARR): A global retention probability calculated from the full dataset. This rate reflects the overall likelihood of a randomly selected customer staying subscribed and is computed using the aggregate churn durations in the dataset. In our case, this is based on each dataset’s customer records available at the time of data collection.
  • Tenure-based Retention Rate (TRR): A personalised retention estimate for each customer, derived from Kaplan-Meier survival curves conditioned on the customer’s current tenure (i.e., account length). This allows us to model retention more granularly, accounting for the fact that long-tenured customers are generally less likely to churn.
@@
The CLV formula uses both ARR and TRR to compute expected profitability. To prevent unrealistically large CLV values due to extremely high retention estimates, we cap the retention rate at a maximum value of 0.995, ensuring numerical stability in the denominator, as discussed in Eq. 3.
3.5 Cost of Intervention
$C_{\text{offer},i}=\text{CPO}\times\text{CLV}_{i}$ (4)
$C_{\text{contact},i}=\max(c_{0},c_{1}\times C_{\text{offer},i})$ (5)
Where:
  • CPO = Cost-per-offer ratio (e.g., 0.1)
  • $c_{0}$ = fixed minimum contact cost
  • $c_{1}$ = proportional multiplier for contact cost
This setup reflects realistic business conditions where retention costs scale with customer value. Complete workflow of e-profits is illustrated in figure 1.
3.6 Comparison with EMP
While the EMP metric also seeks to optimise profit, it differs from e-Profits in several key ways:
  • EMP assumes fixed CLV and global retention parameters [23]; e-Profits uses customer-specific inputs;
  • EMP models offer acceptance with a Beta distribution; e-Profits applies empirically estimated retention probabilities via survival analysis;
  • EMP is analytically optimised for a specific target fraction; e-Profits enables direct computation of total and top-segment profit without restrictive assumptions;
These features make e-Profits more understandable and suitable for operational decision-making in customer-centric businesses.
3.7 Implementation Guidance
e-Profits can be applied to any churn classifier without modifying the model architecture. It requires:
  1. 1.
    Customer-level predictions $\hat{y}_{i}$ and true labels $y_{i}$
  2. 2.
    Monthly revenue and tenure information to compute $\text{CLV}_{i}$
  3. 3.
    Retention rate estimation using Kaplan-Meier survival curves
  4. 4.
    Cost parameters for contact and offers
This modular design makes e-Profits readily integrable into standard evaluation pipelines and dashboards used by marketing and customer analytics teams.
3.8 Toy Example: Manual e-Profits Calculation
To illustrate how e-Profits can help evaluate profitability at the customer level, we present a toy example of five customers with known churn labels, predictions, and financial attributes. Only customers predicted as churners ($\hat{y}_{i}=1$) are considered for intervention, as per Eq. 3.
Table 2: Example customer data for e-Profits illustration
Customer $y_{i}$ (True) $\hat{y}_{i}$ (Pred) $R_{i}$ (Revenue) $r_{i}$ (Retention) Account Length
A 1 1 100 0.80 120
B 0 1 80 0.90 50
C 1 0 120 0.75 180
D 0 0 90 0.85 90
E 1 1 150 0.70 200
[…]
$\text{CLV}_{i}=\frac{R_{i}\cdot M}{1-min(r_{i},0.995)}$
The cost of intervention for each predicted churner is calculated using:
$C_{\text{offer},i}=\text{CPO}\cdot\text{CLV}_{i},\quad C_{\text{contact},i}=\max(c_{0},c_{1}\cdot C_{\text{offer},i})$
Customer A (TP):
$\text{CLV}_{A}=\frac{100\cdot 0.3}{1-0.80}=150,\quad C_{\text{offer}}=15,\quad C_{\text{contact}}=max(5,4.5)$
$\Pi_{A}=150-15-5.0=\mathbf{130}$
Customer B (FP):
$\text{CLV}_{B}=\frac{80\cdot 0.3}{1-0.90}=240,\quad C_{\text{offer}}=24,\quad C_{\text{contact}}=max(5,7.2)$
$\Pi_{B}=-24-7.2=\mathbf{-31.2}$
Customer C and D:
$\hat{y}_{C}=\hat{y}_{D}=0\Rightarrow\Pi_{C}=\Pi_{D}=0$
Customer E (TP):
$\text{CLV}_{E}=\frac{150\cdot 0.3}{1-0.70}=150,\quad C_{\text{offer}}=15,\quad C_{\text{contact}}=max(5,4.5)$
$\Pi_{E}=150-15-5=\mathbf{130}$
Total e-Profits:
$\sum\Pi_{i}=130+(-31.2)+0+0+130=\mathbf{228.8}$
This example illustrates how e-Profits accounts for both the financial value and the prediction correctness of each customer. While both Customers A and E contribute positively, Customer B, although flagged as a churner, incurs a financial loss due to being a false positive. By incorporating retention probability, revenue, and intervention cost, e-Profits provides a more financially meaningful evaluation than traditional classification metrics.
@@ 3.9 Datasets
We apply the e-Profits metric using two publicly available real-world churn datasets that meet the following criteria:
  • Actively used in recent churn prediction research
  • Publicly accessible for reproducibility
  • Moderate classification difficulty (accuracy or F1-score typically <85%)
@@ 3.10 Machine Learning Models
A prior systematic review of churn prediction techniques [28], tree-based ensemble data-driven learning techniques consistently demonstrated superior performance in predictive accuracy and robustness across various datasets. In particular, gradient-boosted techniques such as XGBoost, LightGBM, and CatBoost were frequently reported as top performers in learning accurate models, while Random Forest remained a strong baseline. Moreover, Explainable Boosting Machines were noted for offering a compelling balance between model accuracy and interoperability, an increasingly important consideration in decision-support applications [31]. In this study, we selected the following learning techniques for evaluation:
  • Random Forest (RF): Bagging-based ensemble proved to handle overfitting and high-dimensional data well [32].
  • Gradient Boosting Classifier (GBC): A boosting algorithm that sequentially reduces residual errors [33].
  • XGBoost (XGB): An optimised gradient boosting learning technique known for scalability and handling regularisation well[34].
  • LightGBM (LGBM): A leaf-wise boosting algorithm with histogram-based split search optimal for accelerated training [35].
  • CatBoost (CBC): a technique that can handle categorical variables natively and reduces prediction shift via ordered boosting [36].
  • Explainable Boosting Machine (EBM): A generalised additive model with competitive performance and enhanced interpretability [37].
All models were trained using a 5-fold cross-validation strategy given the datasets do not contain many records. Table 3 shows the hyperparameters and their ranges. This ensured fair performance comparison while accounting for model-specific tuning characteristics.
@@ 3.11 Evaluation Metrics
  • Expected Maximum Profit (EMP): A profit-based metric assuming fixed CLV and offer acceptance probabilities [16].
@@
Metrics are reported in two scenarios: the full population evaluation, which assesses the total profit across all predicted churners, and the top-segment evaluation, focusing on the profit generated from the top 20% of the most profitable customers based on predicted risk. This dual perspective allows us to understand both global model performance and targeted campaign efficiency [^4] Further mathematical formulations for traditional evaluation metrics can be found in our earlier review work [28].
4 Results
This section presents and analyses the performance of several machine learning classifiers for churn prediction using the two benchmark datasets: IBM Telco and Maven Telecom. Each classifier is evaluated using both traditional metrics (e.g., Accuracy, F1-score, AUC) and business-aligned metrics (e.g., Top-Decile Lift, Lift Index, EMP, and the proposed e-Profit. We aim to assess how model rankings and decision quality shift when profitability is considered alongside classification performance.
@@ 4.1 IBM Dataset Analysis
We begin with the IBM Telco dataset, which contains customer tenure and service attributes relevant for churn prediction. Prior to model evaluation, we applied Kaplan-Meier survival analysis to estimate customer retention probabilities. The average retention rate across the IBM dataset is approximately 77%. Figure 2 provides a compact overview, combining the retention curve, lift index, and ROC curve for visual comparison.
(a) Kaplan-Meier estimated retention rate
@@
To assess how well models prioritise high-risk churners, we computed the lift index for the best-performing model under default AUC tuning (EBM). The top-decile lift of 3.01 indicates that the top 10% of customers identified by the model are approximately three times more likely to churn than average, supporting targeted interventions.
Similarly, ROC curves under e-Profits tuning (Figure 2c) show that EBM delivers the best AUC, reaffirming its strong discrimination ability.
@@ 4.1.1 Classifier Performance across Metrics
Tables 4 to 8 report the evaluation outcomes for different models under five tuning objectives: AUC, Accuracy, F1-score, EMP, and e-Profits. For each setting, we report both full-population and top-segment e-Profits, as well as corresponding EMP, Lift, and classification metrics.
Table 4 reports results when classifiers are tuned using AUC. EBM outperforms others in classification metrics such as AUC, Accuracy, and F1-score. However, when ranked by top-segment e-Profits, CatBoost and LGBM exceed EBM. This indicates that AUC-optimised models may not always prioritise high-value customers effectively.
@@
To enrich our analysis of model profitability, we report e-Profits using two variations of retention estimation: the global Average Retention Rate (ARR) and the customer-specific Tenure-based Retention Rate (TRR). Both are derived via Kaplan-Meier survival analysis, as described in the Methodology (see Section 3). This dual reporting allows us to examine how using personalised retention estimates affects model rankings under business-aligned metrics.
Table 4: Classifier performance on IBM dataset when tuned for AUC: Includes both traditional and profit-based evaluation metrics, including e-Profits (ARR/TRR) and EMP under full and top-segment scenarios.
Name F1 (%) Acc. (%) AUC (%) EMP TDL Lift Index e-Prof (0.2, 0.3, ARR) e-Prof (0.2, 0.3, TRR) e-Prof (1.0, 0.3, ARR) e-Prof (1.0, 0.3, TRR) EMP (0.2, 200, 10, 1) EMP (1.0, 200, 10, 1)
LGBM 56.25 78.80 69.82 10.65 2.9659 0.7981 25073 18952 224246 359836 33.35 10.65
RF 57.20 80.03 70.34 10.83 3.0008 0.8075 4532 3606 190975 335472 34.91 10.83
CBC 57.82 79.70 70.77 10.77 2.9833 0.8045 38184 29420 259429 390116 34.41 10.77
GBC 58.56 79.84 71.25 10.84 2.8787 0.8056 8910 7921 225756 374947 34.59 10.84
XGB 59.26 80.22 71.67 10.85 2.9659 0.8068 18365 13631 232821.29 378301 35.20 10.85
EBM 60.06 80.93 72.11 10.95 3.0182 0.8115 19292 15929 253196 399408 36.31 10.95
*CBC = CatBoost, e-prof = e-Profits, Acc. = Accuracy, TDL = Top Decile Lift, ARR = Average Retention Rate, TRR = Tenure Retention Rate
Table 5: Classifier performance on IBM dataset when tuned for Accuracy: Full-spectrum evaluation across classification, lift, and financial metrics, comparing e-Profits and EMP variants.
Name F1 (%) Acc. (%) AUC (%) EMP TDL Lift Index e-Prof (0.2, 0.3, ARR) e-Prof (0.2, 0.3, TRR) e-Prof (1.0, 0.3, ARR) e-Prof (1.0, 0.3, TRR) EMP (0.2, 200, 10, 1) EMP (1.0, 200, 10, 1)
LGBM 56.25 78.80 69.82 10.65 2.9659 0.7981 25073 18952 224246 359836 33.35 10.65
RF 57.20 80.03 0.7034 10.83 3.0008 0.8075 4532 3606 190975 335472 34.91 10.83
XGB 57.88 80.03 70.78 10.86 2.9659 0.8080 16144 12942 229247 372116 34.73 10.86
GBC 58.56 79.84 71.25 10.84 2.8787 0.8056 8910 7921 225756 374947 34.59 10.84
CBC 59.52 80.69 71.78 10.85 3.0008 0.8080 13276 10727 245988 390454 35.41 10.85
EBM 60.06 80.93 72.11 10.95 3.0182 0.8115 19292 15929 253196 399408 36.31 10.95
Table 6: Classifier performance on IBM dataset when tuned for F1-score: Financially-informed metrics reveal shifts in model rankings under e-Profits and EMP.
Name F1 (%) Acc. (%) AUC (%) EMP TDL Lift Index e-Prof (0.2, 0.3, ARR) e-Prof (0.2, 0.3, TRR) e-Prof (1.0, 0.3, ARR) e-Prof (1.0, 0.3, TRR) EMP (0.2, 200, 10, 1) EMP (1.0, 200, 10, 1)
LGBM 56.25 78.80 69.82 10.65 2.9659 0.7981 25073 18952 224246 359836 33.35 10.65
RF 57.20 80.03 70.34 10.83 3.0008 0.8075 4532 3606 190974 335472 34.91 10.83
CBC 57.82 79.70 70.77 10.77 2.9833 0.8045 38184 29420 259429 390116 34.41 10.77
GBC 58.56 79.84 71.25 10.84 2.8787 0.8056 8910 7921 225756 374947 34.59 10.84
XGB 59.26 80.22 71.67 10.85 2.9659 0.8068 18365 13631 232821 378301 35.20 10.85
EBM 60.06 80.93 72.11 10.95 3.0182 0.8115 19292 15929 253196 399408 36.31 10.95
Table 7: Classifier performance on IBM dataset when tuned for EMP: Traditional profit metric shows limitations when compared with e-Profits under personalised cost and value assumptions.
Name F1 (%) Acc. (%) AUC (%) EMP TDL Lift Index e-Prof (0.2, 0.3, ARR) e-Prof (0.2, 0.3, TRR) e-Prof (1.0, 0.3, ARR) e-Prof (1.0, 0.3, TRR) EMP (0.2, 200, 10, 1) EMP (1.0, 200, 10, 1)
XGB 19.78 75.44 55.28 10.91 2.7914 0.8059 0 0 13501 53261 34.41 10.91
GBC 32.60 76.72 59.16 10.65 2.6867 0.7951 0 0 28349 100746 32.34 10.65
CBC 44.58 77.99 63.76 10.76 2.7216 0.8031 0 0 76829 193024 33.62 10.76
RF 54.83 79.18 68.94 10.52 2.8961 0.7932 29209 20091 221862 343511 33.61 10.52
LGBM 56.25 78.80 69.82 10.65 2.9659 0.7981 25073 18952 224246 359836 33.35 10.65
EBM 58.60 80.41 71.20 11.00 2.9833 0.8099 8483 6959 219882 365990 35.20 11.00
Table 8: Classifier performance on IBM dataset when tuned directly for e-Profits: Demonstrates the financial benefit of model selection using customer-specific value, retention, and intervention costs.
Name F1 (%) Acc. (%) AUC (%) EMP TDL Lift Index e-Prof (0.2, 0.3, ARR) e-Prof (0.2, 0.3, TRR) e-Prof (1.0, 0.3, ARR) e-Prof (1.0, 0.3, TRR) EMP (0.2, 200, 10, 1) EMP (1.0, 200, 10, 1)
RF 54.10 78.56 68.51 10.42 2.8438 0.7918 20665 13917 221891 345815 32.05 10.42
LGBM 56.25 78.80 69.82 10.65 2.9659 0.7981 25073 18952 224246 359836 33.35 10.65
XGB 56.46 78.47 69.98 10.56 2.9135 0.7939 37805 27632 241516 372267 32.40 10.56
GBC 56.53 78.89 70.00 10.60 2.7565 0.7936 38196 29315 255144 382767 32.19 10.60
CBC 57.82 79.70 70.77 10.77 2.9833 0.8045 38184 29420 259429 390116 34.41 10.77
EBM 60.80 81.02 72.61 10.95 3.0182 0.8108 19445 15953 265068 413470 35.84 10.95
These results underscore that models optimised for ranking correctness may misallocate retention resources in real-world settings where profitability is paramount.
Tables 5 and 6 show model performance when tuned for Accuracy and F1-score, respectively. EBM again dominates classification performance, but the ranking for top-segment e-Profits shifts, with LGBM and CatBoost producing higher financial returns. This demonstrates that models selected using traditional metrics may misalign with business goals, especially in targeted retention campaigns.
These findings strengthen the case for evaluating classifiers by correctness and financial impact, especially in resource-constrained retention settings.
Table 7 provides results when models are tuned using EMP. While EMP reflects a profit orientation, its reliance on fixed cost and acceptance rate assumptions limits its alignment with individual-level profit signals. For instance, despite lower EMP, Random Forest delivers substantially better top-segment e-Profits than CatBoost. This inconsistency shows that even profit-based metrics like EMP can miss context-specific opportunities.
Notably, several models show zero profit in the top segment under EMP tuning, an artefact of global assumptions about CLV and offer response, not model capability.
Table 8 shows results when models are tuned directly on our proposed e-Profits metric. While EBM achieves the highest overall profit, GBC, CatBoost, and XGBoost yield superior outcomes for the top 20% most profitable customers. This is also illustrated by radar graph in figure 3. This divergence suggests that e-Profits is better suited for applications where ROI is maximised by targeting smaller, high-value segments.
By enabling per-customer evaluation using actual retention behaviour, cost, and value, e-Profits provides a practical basis for real-world model selection.
4.1.2 Key Insights
  • Traditional metrics (AUC, F1) fail to consistently rank models by business value.
  • EMP improves this alignment but remains sensitive to fixed parameter assumptions.
  • e-Profits offers a more faithful and granular evaluation by incorporating per-customer economics.
  • Classifier rankings change substantially when moving from AUC to e-Profits optimisation, highlighting the risk of metric misalignment in real-world deployment.
Figure 3: Radar graph for IBM dataset (Tuned on e-profits considering 20% profitable customers and tenure based retention)

### v2

@@ 3.2.4 Retention Probability Estimation via Survival Analysis
  • Tenure-based Retention Probability (TRR): A customer-level retention estimate that varies with the customer’s current tenure $T_{i}$, computed from a population Kaplan-Meier survival curve. Let $T$ denote the customer lifetime random variable and let $S(t)$ be the Kaplan-Meier survival function. For a customer $i$ with current tenure $T_{i}$, we define the TRR as the one-period conditional survival probability over horizon $\Delta$:
    $r_{i}^{\text{TRR}}(\Delta)=\Pr(T>T_{i}+\Delta\mid T>T_{i})=\frac{S(T_{i}+\Delta)}{S(T_{i})},\qquad i=1,\dots,N.$ (6)
    In our experiments we set $\Delta=1$ month to align retention with the monthly revenue definition used in Eq. 3. This definition explicitly conditions on the customer having already survived up to tenure $T_{i}$, and therefore matches the intended meaning of conditioned on current tenure.
  • Average Retention Probability (ARR): The average retention Probability is defined as the mean conditional survival probability over the empirical tenure distribution:
    $r^{\text{ARR}}(\Delta)=\frac{1}{N}\sum_{i=1}^{N}r_{i}^{\text{TRR}}(\Delta)=\frac{1}{N}\sum_{i=1}^{N}\frac{S(T_{i}+\Delta)}{S(T_{i})}.$ (7)
    Equivalently, in continuous form:
    $r^{\text{ARR}}(\Delta)=\int_{0}^{\infty}\frac{S(t+\Delta)}{S(t)}\,dF(t),$ (8)
@@
where $F(t)$ is the empirical distribution of observed customer tenures. The CLV formula uses both ARR and TRR to compute expected profitability. To prevent unrealistically large CLV values due to extremely high retention estimates, we cap the retention probability at a maximum value of 0.995, ensuring numerical stability in the denominator, as discussed in Eq. 3. Kaplan-Meier survival curves used to compute retention inputs (TRR/ARR) were fitted using only the training split. The held-out test split was never used to fit the survival curve or compute retention.
During hyperparameter selection, we perform 5-fold cross-validation on the training split only. For each tuning target (AUC, Accuracy, F1, EMP, or e-Profits), the corresponding metric is computed on validation folds using retention probabilities derived from a Kaplan-Meier survival curve $S(\cdot)$ that is estimated once on the training split and kept fixed throughout tuning. The Kaplan-Meier curve is therefore not refitted per fold. After hyperparameter selection, the chosen model is refit on the full training split and evaluated once on the held-out test split, which is never used to estimate retention.
Algorithm 1 Retention-probability computation from a Kaplan-Meier curve
1:Fit Kaplan-Meier survival curve $S(\cdot)$ using training churn times and censoring indicators
2:for each customer $i$ do
3:  Observe current tenure $T_{i}$
4:  Compute $r_{i}^{\text{TRR}}(\Delta)\leftarrow\frac{S(T_{i}+\Delta)}{S(T_{i})}$
5:end for
6:Compute $r^{\text{ARR}}(\Delta)\leftarrow\frac{1}{N}\sum_{i=1}^{N}r_{i}^{\text{TRR}}(\Delta)$
3.3 Comparison with EMP and Implementation Guidance
While EMP also aims to optimise profit, it differs from e-Profits in several important respects. EMP assumes fixed customer lifetime value and global retention parameters, whereas e-Profits operates at the customer level. EMP models offer acceptance via a parametric Beta distribution, while e-Profits relies on empirically estimated retention probabilities obtained through survival analysis. Finally, EMP is analytically optimised for a specific target fraction, whereas e-Profits enables direct computation of both total and segment-level profit without imposing restrictive population-level assumptions. These differences make e-Profits more transparent and better suited for operational decision-making. These differences are summarised empirically in Table 3 and 4.
The e-Profits metric can be applied to any churn classifier without modifying the model architecture. Its implementation requires predicted churn labels, customer-level revenue and tenure information to compute lifetime value, retention probability estimation via Kaplan-Meier survival analysis, and specification of contact and offer cost parameters. This modular design allows e-Profits to integrate seamlessly into standard evaluation pipelines and business dashboards used by marketing and customer analytics teams.
3.4 Computational overhead of Kaplan–Meier
Kaplan-Meier estimation requires sorting observed churn/censoring times and computing a stepwise survival curve, with computational cost on the order of $O(n\log n)$ due to sorting. The Kaplan-Meier curve is estimated once using the training split only (to avoid leakage) and then reused to derive retention inputs (ARR/TRR) throughout hyperparameter tuning and test evaluation; hence, KM estimation adds only a small, one-time preprocessing cost. During tuning, the dominant runtime cost arises from repeatedly fitting models and evaluating fold scores across hyperparameter configurations.
Empirically, on the Maven dataset (mean over models), tuning on AUC required 147.48 s, while tuning on e-Profits required 150.12 s (ARR) and 103.50 s (TRR); EMP tuning required 287.50 s. These timings (measured on the Maven dataset) indicate that incorporating Kaplan-Meier derived retention into e-Profits does not add prohibitive overhead relative to standard objectives; the overall runtime is primarily governed by model training and the size of the hyperparameter grid.
@@ 3.5 Datasets
We apply the e-Profits metric using two publicly available real-world telecom churn datasets selected for their widespread use in recent research, public accessibility for reproducibility, and moderate classification difficulty (accuracy or F1-score typically <85%).
@@ 3.6 Machine Learning Models
A prior systematic review of churn prediction techniques [manzoor2024review], tree-based ensemble data-driven learning techniques have consistently demonstrated superior performance in predictive accuracy and robustness across various datasets. In particular, gradient-boosted techniques such as XGBoost, LightGBM, and CatBoost were frequently reported as top performers in learning accurate models, while Random Forest remained a strong baseline. Moreover, Explainable Boosting Machines were noted for offering a compelling balance between model accuracy and interpretability, an increasingly important consideration in decision-support applications [liu2023concrete]. In this study, we selected the following learning techniques for evaluation:
Random Forest (RF): Bagging-based ensemble proved to handle overfitting and high-dimensional data well [bentejac2021comparative].
Gradient Boosting Classifier (GBC): A boosting algorithm that sequentially reduces residual errors [alshourbaji2023efficient].
XGBoost (XGB): An optimised gradient boosting learning technique known for scalability and handling regularisation well [paliwal2022xgbrs].
LightGBM (LGBM): A leaf-wise boosting algorithm with histogram-based split search optimal for accelerated training [sina2022model].
CatBoost (CBC): a technique that can handle categorical variables natively and reduces prediction shift via ordered boosting [asif2025data].
Explainable Boosting Machine (EBM): A generalised additive model with competitive performance and enhanced interpretability [amekoe2024exploring].
All models were trained using a 5-fold cross-validation strategy since the datasets do not contain many records. Table 2 shows the hyperparameters and their ranges. This ensured fair performance comparison while accounting for model-specific tuning characteristics.
@@ 3.7 Evaluation Metrics
Expected Maximum Profit (EMP): A profit-based metric assuming fixed CLV and offer acceptance probabilities [verbraken2012novel].
@@
Metrics are reported in two scenarios: the full population evaluation, which assesses the total profit across all predicted churners, and the top-segment evaluation, focusing on the profit generated from the top 20% of the most profitable customers based on predicted risk. This dual perspective allows us to understand both global model performance and targeted campaign efficiency [^3] Further mathematical formulations for traditional evaluation metrics can be found in our earlier review work [manzoor2024review].
From a deployment perspective, the e-Profits metric is practically employable within firms. It integrates seamlessly with existing modelling workflows and provides direct financial interpretation that supports resource allocation decisions in marketing and sales. This makes it a high-utility tool for internal stakeholders responsible for campaign design and retention optimization.
3.8 Model Selection and Validation Procedure
We split the dataset once into a training set and a held-out test set. Hyperparameters are selected using 5-fold cross-validation on the training set only. For each tuning target (AUC, accuracy, F1, EMP, or e-Profits), the corresponding metric is used as the cross-validation score for grid search and to select the best hyperparameters; the final model is then refit on the full training set using the selected configuration. All reported results provided in Tables 3 and 4 are computed on the held-out test set, which is not used during tuning or model selection.
4 Experimental Results
This section presents and analyses the performance of several machine learning classifiers for churn prediction using the two benchmark datasets: IBM Telco and Maven Telecom. Each classifier is evaluated using both traditional metrics (e.g., Accuracy, F1-score, AUC) and business-aligned metrics (e.g., top-decile lift, Lift Index, EMP, and the proposed e-Profits). We aim to assess how model rankings and decision quality shift when profitability is considered alongside classification performance.
@@ 4.1 IBM Dataset Analysis
We begin with the IBM Telco dataset, which contains customer tenure and service attributes relevant for churn prediction. Prior to model evaluation, we applied Kaplan-Meier survival analysis to estimate customer retention probabilities. The ARR estimated from the training split of the IBM dataset is approximately 0.77. This retention estimate is then treated as fixed and applied when evaluating models on the held-out test split. Figure 2 provides a compact overview, combining the retention curve, lift index, and ROC curve for visual comparison.
(a) Kaplan-Meier estimated retention curve
@@
To assess how well models prioritise high-risk churners, we computed the lift index for the best-performing model under default AUC tuning (EBM). The top-decile lift of 3.01 indicates that the top 10% of customers identified by the model are approximately three times more likely to churn than average, supporting targeted interventions. Similarly, ROC curves under e-Profits tuning (Figure 2c) show that EBM delivers the best AUC, reaffirming its strong discrimination ability.
@@ 4.1.1 Classifier Performance across Metrics
Table 3 reports the evaluation outcomes for different models under five tuning objectives: AUC, Accuracy, F1-score, EMP, and e-Profits, respectively. For each setting, we report both full-population and top-segment e-Profits, as well as corresponding EMP, Lift, and classification metrics.
@@
To enrich our analysis of model profitability, we report e-Profits for both the full population and the top 20% most profitable customers, under two retention assumptions: the global Average Retention Rate (ARR) and the customer-level Tenure-based Retention Rate (TRR). Both are derived via Kaplan-Meier survival analysis, as described in the Methodology (see Section 3). This dual reporting allows us to examine how tenure-conditioned retention estimates affect model rankings under business-aligned metrics.
When optimised on AUC, EBM outperforms other classifiers on traditional predictive metrics, including AUC, Accuracy, F1-score, Top-decile lift, and EMP. However, Random Forest achieves higher e-Profits on both the full-population and the top 20% segment, indicating that AUC-optimised models do not necessarily prioritise customers with the highest financial value. While EBM consistently dominates classification performance across tuning objectives, models such as XGB and LGBM frequently achieve higher financial returns when evaluated using e-Profits.
Examining profit outcomes in more detail reveals systematic differences between full-population and targeted evaluation. LGBM achieves the highest top-segment (20%) e-Profits under both ARR and TRR when tuned on F1-score, EMP, and e-Profits. In contrast, for full-population profitability, XGB yields the highest e-Profits when tuned on Accuracy, F1-score, and e-Profits. When tuned on EMP, EBM attains a slightly higher full-population e-Profits than XGB, with XGB ranked second. These results show that profit-based model rankings differ substantially from those induced by traditional predictive objectives. Crucially, models tuned directly on e-Profits achieve the highest or near-highest realised profit across both full-population and top-segment evaluations, while maintaining competitive predictive performance.
Although EMP reflects a profit-oriented objective, its reliance on fixed cost and acceptance-rate assumptions limits its alignment with individual-level profit signals. As a result, models tuned using EMP do not consistently maximise e-Profits; in particular, several models yield zero profit in the top segment under EMP tuning, an artefact of global CLV and offer-response assumptions rather than model capability. When models are tuned directly on e-Profits, XGB achieves the highest overall profit, while LGBM and CBC yield superior outcomes for the top 20% most profitable customers, as also illustrated by the radar plot in Figure 4.
Table 3: Classifiers’ performance on the IBM dataset under AUC, Accuracy, F1-score, EMP and e-Profits, respectively.
Name F1 (%) Acc. (%) AUC (%) TDL Lift Index e-Prof (0.2, 0.3, ARR) e-Prof (0.2, 0.3, TRR) e-Prof (1.0, 0.3, ARR) e-Prof (1.0, 0.3, TRR) EMP (0.2, 200, 10, 1) EMP (1.0, 200, 10, 1)
Tuned on AUC
RF 57.20 79.18 83.77 2.8961 0.7932 878821 744024 7289696 9104621 33.61 10.52
GBC 57.82 76.72 83.82 2.6867 0.7951 0 0 852922 448270 32.34 10.65
XGB 59.26 75.20 85.62 2.7216 0.8064 0 0 330130 151294 34.57 10.97
CBC 57.82 77.99 84.85 2.7216 0.8031 0 0 2311538 1957368 33.62 10.76
LGBM 56.25 79.03 84.72 2.9659 0.8003 0 0 2150792 7944468 33.61 10.67
EBM 59.62 80.83 86.16 3.0008 0.8103 478040 701949 6675152 7894896 35.89 10.98
Tuned on Accuracy
RF 57.20 80.03 85.64 3.0008 0.8075 136351 199609 5745843 6891397 34.91 10.83
GBC 58.56 79.84 85.38 2.8787 0.8056 268062 511788 6792305 8518455 34.59 10.84
XGB 59.32 80.27 85.48 2.8612 0.8051 608452 922701 7787442 9953379 34.57 10.87
CBC 59.52 80.69 85.74 3.0008 0.8080 399446 614876 7059167 9283514 35.41 10.85
LGBM 55.04 79.51 85.25 3.0182 0.8038 574879 855865 5741153 6941722 33.55 10.74
EBM 60.61 81.12 86.15 3.0182 0.8118 565626 823097 7401027 8898532 36.15 10.93
Tuned on F1-score
RF 57.20 80.03 85.64 3.0008 0.8075 136351 199609 5745843 6891397 34.91 10.83
GBC 58.56 79.84 85.38 2.8787 0.8056 268062 511788 6792305 8518455 34.59 10.84
XGB 59.51 80.36 85.58 2.9310 0.8071 668892 894586 7805429 9953379 34.96 10.86
CBC 57.82 79.70 85.21 2.9833 0.8045 754358 1060291 7365250 9193007 34.41 10.77
LGBM 56.25 78.80 84.25 2.9659 0.7981 1148846 1583632 6746884 8421249 33.35 10.65
EBM 60.61 81.12 86.15 3.0182 0.8118 608452 922701 7787442 9833881 36.15 10.93
Tuned on EMP
RF 51.11 79.08 85.43 2.8438 0.8052 0 0 322641 3204003 34.43 10.81
GBC 56.09 79.70 85.27 2.8787 0.8051 10062 81201 5411529 6497417 34.09 10.79
XGB 59.52 80.69 85.74 3.0008 0.8080 399446 614876 7701027 9783514 35.41 10.85
CBC 57.67 80.27 85.65 3.0008 0.8085 136351 199609 5793988 6956820 35.04 10.85
LGBM 56.87 79.60 84.68 3.0531 0.7998 686595 1069797 5784574 7082188 33.93 10.64
EBM 60.84 81.12 86.22 2.9833 0.8124 573412 889289 7732943 9829970 35.70 10.95
Tuned on e-Profits (under TRR on full population)
RF 54.10 78.56 83.47 2.8438 0.7918 621741 584309 8776019 9150798 32.05 10.42
GBC 56.53 78.89 83.73 2.7565 0.7936 1149186 1426809 7676514 9632075 32.19 10.60
XGB 58.47 79.22 83.82 2.8612 0.7937 1014112 1240208 8312032 10991460 32.69 10.61
CBC 57.82 79.70 85.21 2.9833 0.8045 1148846 1583632 7805429 9933881 34.41 10.77
LGBM 57.09 78.66 83.50 2.8961 0.7913 1571624 1627728 7530726 9564945 32.74 10.48
EBM 59.05 80.50 86.09 2.966 0.8110 552421 851399 7339760 9785358 36.05 10.94
*e-Profits are reported as $e\text{-}Prof(\eta,M,\mathrm{RR})$, where $\eta\in\{0.2,1.0\}$ denotes the evaluated customer fraction (top 20% or full population), $M=0.3$ is the profit margin used in Eq. 3, and $\mathrm{RR}\in\{\mathrm{ARR},\mathrm{TRR}\}$ specifies whether average or tenure-based retention probabilities are used. EMP values are computed as $\mathrm{EMP}(\eta,\mathrm{CLV},C_{\text{offer}},C_{\text{contact}})$ with $\mathrm{CLV}=200$, $C_{\text{offer}}=10$, and $C_{\text{contact}}=1$.
*CBC=CatBoost, Acc.=Accuracy, TDL=Top-Decile Lift